OpenStudy (anonymous):

Tricky integral: \[\int e^{-2x}\left[(e^x)^2 +(x^x)^2\right]\ln x dx\] No Googling allowed :)

OpenStudy (solomonzelman):

u=e^x

OpenStudy (anonymous):

haha no googling allowed,what about wofram????

OpenStudy (solomonzelman):

I mean u=e^2x

OpenStudy (anonymous):

Go right ahead @Jonask, just stay away from Math Stack Exchange!

OpenStudy (solomonzelman):

\[\int\limits_{ }^{ }e^{-2x}2e^{2x}~dx\]\[\int\limits_{ }^{ }2~dx\]and it is 2x?

OpenStudy (solomonzelman):

Ohh the ln. sorry

OpenStudy (solomonzelman):

it would be integral of 2ln(x) I guess. After simplifying the Es

OpenStudy (anonymous):

\[\int \ln x dx+\int \frac{x^{2x}}{e^{2x}}\ln xdx\]

OpenStudy (anonymous):

@SolomonZelman your simplification isn't right. The first term should simplify to 1 (multiplied by \(\ln x\)).

OpenStudy (solomonzelman):

but in your brackets isn't it just 2e^2x ?

hartnn (hartnn):

i see the derivative of x^x in the function

OpenStudy (solomonzelman):

and e^2x * e^-2x= e^(-2x + 2x) = e^0 = 1 So, it would be with a 2.

hartnn (hartnn):

which is x^x (1+ln x)

OpenStudy (anonymous):

Second term is \(\large x^x\), not \(\large e^x\)... Yes, @hartnn is on the right track. It's a matter of coming up with the right sub now.

OpenStudy (solomonzelman):

ohh.. I am blind again

ganeshie8 (ganeshie8):

\[\large \int e^{-2x}\left[(e^x)^2 +(x^x)^2\right]\ln x dx = \int (1 + e^{2x(\ln x - 1)}) \ln x \] sub 2x(lnx-1) = u ?

OpenStudy (anonymous):

Not what I would have suggested @ganeshie8, but it might still lead to a solution.

hartnn (hartnn):

first integral is ln x u = e^(-2x) x^(2x) for the 2nd integral

hartnn (hartnn):

du = 2 e^(-2x) x^(2x) ln x

hartnn (hartnn):

*dx

hartnn (hartnn):

\(\int \ln x dx+\int \frac{x^{2x}}{e^{2x}}\ln xdx \\u = \frac{x^{2x}}{e^{2x}} \\ du =2\frac{x^{2x}}{e^{2x}} dx \\\int \ln x dx+\int \frac{du}{2}\) and the rest is history :P

ganeshie8 (ganeshie8):

Clever! :)

OpenStudy (anonymous):

Nice work! Honorable mention @Jonask, you had the right approach as well

hartnn (hartnn):

was just trying u =x^(2x) before! then realized how simple u = e^(-2x)x^(2x) made this problem :D

hartnn (hartnn):

thanks for the good question! :)

OpenStudy (anonymous):

i deleted i thought it was not gonna work

OpenStudy (anonymous):

the rest is history,nice! @hartnn

hartnn (hartnn):

lol, too lazy to write out those easy steps :P

hartnn (hartnn):

bdw, integral of ln x would be solved using integration by parts

hartnn (hartnn):

*** \(\large du =2\frac{x^{2x}}{e^{2x}} \ln xdx\)